# Quantum Monte Carlo and Stabilizer States

### v1

@@ Abstract
The Quantum-Monte-Carlo technique known as the Stochastic Series Expansion (SSE) relies on a crucial no-branching condition: the SSE sampling is carried out in the computational basis, and the no-branching assumption ensures that superpositions of basis-states do not appear when operators are applied. Without this proviso, the number of complex amplitudes would grow exponentially with the number of qubits and would eventually overwhelm the memory and processing power of a classical computer. However, the action of Clifford group elements on stabilizer states can be very efficiently described without resorting to an amplitude description. We explore how stabilizer states allow an extension of the SSE technique, and we give an example of a toy model that can be studied in this way.
@@ I Introduction
The difficulty of classically simulating quantum systems stems from the typically enormous Hilbert spaces and the need to sum complex probability amplitudes over all possible paths from the initial to the final state. The latter summation can lead to delicate cancellations that can be difficult to keep track of. Nevertheless, over the past decades, clever computational tools have been developed for certain physical systems. Examples of these are the various types of Quantum Monte Carlo simulations based on Euclidean path integrals Hirsch et al. (1982); Beard and Wiese (1996); Prokof’Ev et al. (1998). They all rely on the fact that occasionally it is possible to rewrite a system’s thermal or ground state in terms of purely real and non-negative probability amplitudes. Then the latter can be interpreted as weights of states that can be sampled using the Markov chain Monte Carlo approach Metropolis et al. (1953). However, this approach can fail when the Hamiltonian has a sign problem that persists under local basis change (i.e. after acting with a low-depth unitary circuit). Usually, this is interpreted as a case where quantum mechanics is an essential component of the problem and obstructs efficient classical sampling Li and Yao (2019); Henelius and Sandvik (2000). Contrast this with the situation in quantum computing, where it was shown by Lloyd that a universal quantum system can simulate any other quantum system efficiently Lloyd (1996), proving a statement that was conjectured by Feynman in Feynman (2018).
@@
Despite much work, the limits of classical simulability are not well understood. Further progress may be made by pushing the boundaries on what is classically simulable. In this paper, we contribute to this discussion by exploring an extension of the stochastic series expansion (SSE) Monte Carlo technique Sandvik (2010); Sandvik and Kurkijärvi (1991); Sandvik (1999); Kaul et al. (2013). The SSE technique relies on a high-temperature expansion of the partition function. When the expansion terms can be interpreted as non-negative weights, a Monte Carlo sampling is often possible. During the simulation, powers of the Hamiltonian act on quantum states. In order to make this tractable on a classical computer, the SSE technique uses a no-branching provision that avoids the formation of quantum superpositions. As a consequence of this, it is not possible to arbitrarily modify an operator string by inserting or removing an (off-diagonal) operator, since this will quickly result in a zero-weight state. Thus, inspired by the algorithm of Swendsen and Wang Swendsen and Wang (1987), many clever loop and cluster algorithms have been proposed over the years to sample the state and operator configurations efficiently Sandvik (1999, 2003); Syljuåsen and Sandvik (2002).
@@
The present paper shows that the no-branching condition can be removed even on a classical computer for some models by making use of stabilizer states. These states have been extensively used in the construction of quantum stabilizer codes Shor (1995); Calderbank and Shor (1996); Steane (1996); Gottesman (1997). They are distinguished within Hilbert space by having an elegant formalism for their description that enables rapid manipulation under the action of the Clifford group. Despite their classical tractability, they can have high degrees of quantum entanglement Fattal et al. (2004).
@@
We start in section II with a brief summary of the stochastic series expansion and introduce the idea of stabilizer states and their compact description. In section III we demonstrate the use of stabilizers states in Quantum Monte Carlo simulations by picking a toy Hamiltonian and illustrating the algorithm in detail. The GitHub repository Jeevanesan contains the C++ implementation. Finally, section IV sketches how to extend this method to the transverse-field Ising model and certain $\mathbb{Z}_{2}$ gauge theories.
@@ II A brief review of SSE Quantum Monte Carlo and the Stabilizer Formalism
Consider a Hamiltonian on $N$ qubits. In a first step, the partition function is expanded in powers of $\beta$
$Z=\text{Tr}[e^{-\beta H}]=\sum_{s}\langle s|e^{-\beta H}|s\rangle=\sum_{s}\sum_{n=0}^{\infty}\frac{(-\beta)^{n}}{n!}\langle s|H^{n}|s\rangle$ (1)
@@
$Z=\sum_{s}\sum_{n=0}^{\infty}\frac{(-\beta)^{n}}{n!}\langle s|H_{i_{1}}H_{i_{2}}\dots H_{i_{n}}|s\rangle$ (2)
@@
where the factor (${{L}\choose{L-n}}^{-1}$ compensates for overcounting identical terms. The sum is extended over all combinations of length-$L$ operator strings and $n$ refers to the number of non-identity operators in it. In the second line we introduced the notation $W$ for the weight of each configuration.
@@
To summarize, the partition function (1) has been transformed into a sum of matrix elements involving $L$ operators. Each term can be interpreted as the weight of a physical process where the system begins and ends in state $|s\rangle$ after the action of $L$ operators. Thus, if the Hamiltonian $H$ describes a lattice system of dimension $d$ then the representation (4) describes a system of dimension $d+1$ where the additional dimension is periodic with $L$ lattice sites Suzuki (1976).
@@
The crucial step in SSE Monte Carlo is that the decomposition of $H$ into terms $\sum_{i}H_{i}$ is done in such a way that the action of any $H_{i}$ on a computational basis state $|s\rangle$ yields another computational basis state $|s^{\prime}\rangle$. This can always be accomplished by breaking $H$ into sufficiently small pieces and is called the no-branching provision. It guarantees that no superpositions over basis states are formed during the simulation process and thereby ensures easy bookkeeping. General superpositions of basis states are practically impossible to keep track of on a classical computer, except when system sizes are small.
@@
The sampling of the partition function now proceeds by setting up a Markov chain Metropolis et al. (1953); Diaconis (2009) with each element of the chain being a state $|s\rangle$ together with a string of operators. The weight of the Markov chain element is given by the corresponding term in eq. (3). An efficient simulation is only possible if the weights are all non-negative. Occasionally, one can manipulate the Hamiltonian terms $H_{i}$ by the addition of constants and render all weights non-negative. If this is not possible, one faces the Monte-Carlo sign problem. It can be dealt with, in a brute force manner, by absorbing the sign of the weight into the observable. But generally, this leads to long convergence times when calculating observable averages, and the variances diverge exponentially with system-size.
@@
A Markov chain update occurs by proposing a change of the state $|s\rangle$ or the insertion/deletion of an operator in the string. The proposal is accepted or rejected to satisfy detailed balance. The latter involves the ratios of the weights, an example appears in section III.2. Note that if only one operator is inserted or removed, most factors in the weights are unchanged and will therefore cancel out. This leads to simple expressions for the update rules. However, since the application of the $H_{i_{k}}$ in eq. (3) always yields basis states, randomly inserting or rejecting an operator will often result in a state with zero weight. Instead, the insertion and removal of operators have to be carefully crafted. For certain Hamiltonians, smart algorithms have been discovered that insert/delete a large number of operators simultaneously in order to achieve fast changes of the sampled configurations Syljuåsen and Sandvik (2002); Sandvik (2003, 1999).
@@
In this paper, we ask whether the no-branching condition is essential or if there is an efficient way to track the branching of the states. To this end, we return to the expression in eq. (3) and consider an interpretation of the sequence of operators as a quantum circuit. This is possible if each $H_{i}$ is either proportional to a unitary or to a projection operator, with the latter describing the outcomes of partial projective measurements. This observation by itself would not be particularly useful since an arbitrary quantum circuit still takes an exponential time to simulate classically. However, if the unitary is a Clifford circuit, efficient simulation is possible, as codified in the theorem of Gottesman and Knill Gottesman (1998a).
@@
Thus, applying $U$ to $|\psi\rangle$ is equivalent to conjugating the generators by $U$.
For general $U$, the resulting generators will lie outside the Pauli group and their description will be just as cumbersome as describing $|\psi\rangle$ in the computational basis. However, if $U$ maps all Pauli group elements to (possibly different) Pauli group elements under conjugation, then an update according to (9) is straightforward. Since conjugation by $U$ followed by conjugation by $V$ is the same as the joint conjugation by $VU$, these elements form a group, the Clifford group. It was shown in Gottesman (1998b) that the Clifford group is finitely generated by the Hadamard, phase and controlled-not gates. This set of gates is, of course, not universal, which is why circuits made up of only these elements are quickly simulable with classical algorithms. Including $T$ gates would make the circuit universal and generally classically intractable.
@@ III A Controlled-Not Toy Model
The two sums do not commute with each other. The CX term is a controlled-not operation with qubit $i$ being the control and qubit $i+1$ the target. In the Z-basis, this operator flips qubit $i+1$, conditioned on the state of qubit $i$ being $|1\rangle$. We work throughout this paper with periodic boundary conditions such that site $N+1$ is identified with site $1$. We also measure energy in units of $J$, i.e. we set $J=1$. Conditional interactions, similar to the CX operator, also appear in constrained statistical mechanics systems and lead to interesting dynamical effects. A recent example is the PXP model that describes Rydberg blockade physics Lesanovsky and Katsura (2012): an atom can only be excited if the neighboring atoms are in the ground state. The system shows interesting many-body dynamics Bernien et al. (2017) and quantum scarring Serbyn et al. (2021) that can be traced back to the constrained dynamics.
It is possible to write the Hamiltonian in eq. (10) because CX is not only unitary but also hermitian. Thus, in terms of Pauli operators, the Hamiltonian eq. (10) reads
@@
This simple example already illustrates how the consecutive application of the sequence of operators in eq. (15) results in superpositions of computational basis states with non-zero entanglement. The fact that the Bell-state appears in eq. (18) also demonstrates that in general these superpositions cannot be rewritten as product states by changing basis. Nevertheless, the stabilizer formalism allows us to keep track of such states.
@@ III.1 Calculation of $\mathcal{M}$
$G_{m}$$\equiv$ $\gamma_{m}X_{1}^{P_{m1}}\otimes\dots\otimes X_{N}^{P_{mN}}\otimes Z_{1}^{Q_{m1}}\otimes\dots\otimes Z_{N}^{Q_{mN}}$ (19)
$=$ $\gamma_{m}\prod_{n=1}^{N}X_{n}^{P_{mn}}\prod_{n=1}^{N}Z_{n}^{Q_{mn}}$
@@
This step is necessary because the stabilizers do not fix the norm of a state. After projector $\Pi_{i}$ acts on state $|\psi\rangle$, the new stabilizers can be worked out as follows. Since the generators are products of Pauli operators, the $X_{i}$ operator will either commute or anti-commute with them. Retain all the generators that commute with $X_{i}$. If all generators commute with $X_{i}$, we are done and set $F^{\prime}=F$.
@@
$\gamma^{\prime}_{m_{a}}$ $=$ $\gamma_{m_{1}}\gamma_{m_{a}}$ (31)
@@
All other components of $\gamma$, $P$ and $Q$ are left unmodified.
@@
We first observe that $\langle s|s^{\prime}\rangle$ is non-negative. The reason for this is that both $\Pi_{i}$ and $[\text{CX}]_{i,i+1}$ are non-negative matrices in the $Z$ basis, i.e. all the entries are non-negative. The product of non-negative matrices is non-negative in that same basis, thus $\langle s|s^{\prime}\rangle$ is non-negative.

### v2

@@ Abstract
The Quantum Monte Carlo technique known as the Stochastic Series Expansion (SSE) relies on a crucial no-branching condition: the SSE sampling is carried out in the computational basis, and the no-branching assumption ensures that superpositions of basis-states do not appear when operators are applied. Without this proviso, the number of complex amplitudes would grow exponentially with the number of qubits and would eventually overwhelm the memory and processing power of a classical computer. However, the action of Clifford group elements on stabilizer states can be very efficiently described without resorting to an amplitude description. We explore how stabilizer states allow an extension of the SSE technique, and we give an example of a toy model that can be studied in this way.
@@ I Introduction
The difficulty of classically simulating quantum systems stems from the typically enormous Hilbert spaces and the need to sum complex probability amplitudes over all possible paths from the initial to the final state. The latter summation can lead to delicate cancellations that can be difficult to keep track of. Nevertheless, over the past decades, clever computational tools have been developed for certain physical systems. Examples of these are the various types of Quantum Monte Carlo simulations based on Euclidean path integrals Hirsch et al. (1982); Beard and Wiese (1996); Prokof’Ev et al. (1998). They all rely on the fact that occasionally it is possible to rewrite a system’s thermal or ground state in terms of purely real and non-negative probability amplitudes. Then the latter can be interpreted as weights of states that can be sampled using the Markov chain Monte Carlo approach Metropolis et al. (1953). However, this approach can fail when the Hamiltonian has a sign problem that persists under local basis change (i.e. after acting with a low-depth unitary circuit). Usually, this is interpreted as a case where quantum mechanics is an essential component of the problem and obstructs efficient classical sampling Li and Yao (2019); Henelius and Sandvik (2000). Contrast this with the situation in quantum computing, where it was shown by Lloyd that a universal quantum system can simulate any other local quantum system efficiently Lloyd (1996), proving a statement that was conjectured by Feynman in Feynman (2018).
@@
Despite much work, the limits of classical simulability are not well understood. Further progress may be made by pushing the boundaries on what is classically simulable. In this paper, we contribute to this discussion by exploring an extension of the stochastic series expansion (SSE) Monte Carlo technique Sandvik (2010); Sandvik and Kurkijärvi (1991); Sandvik (1999); Kaul et al. (2013). The SSE technique relies on a high-temperature expansion of the partition function. When the expansion terms can be interpreted as non-negative weights, a Monte Carlo sampling is often possible. During the simulation, powers of the Hamiltonian act on quantum states. In order to make this tractable on a classical computer, the SSE technique uses a no-branching provision that avoids the formation of quantum superpositions. As a consequence of this, it is not possible to arbitrarily modify an operator string by inserting or removing an (off-diagonal) operator, since this will quickly result in a zero-weight state. Thus, inspired by the algorithm of Swendsen and Wang Swendsen and Wang (1987), many clever loop and cluster algorithms have been proposed over the years to sample the state and operator configurations efficiently Sandvik (1999, 2003); Syljuåsen and Sandvik (2002). The SSE technique has also been extended to quantum computers in form of a quantum SSE algorithm Tan et al. (2022), see also Jiang et al. (2024) for a review.
@@
The present paper shows that the no-branching condition can be removed even on a classical computer for some models by making use of stabilizer states. These states have been extensively used in the construction of quantum stabilizer codes Shor (1995); Calderbank and Shor (1996); Steane (1996); Gottesman (1997). They are distinguished within Hilbert space by having an elegant formalism for their description that enables rapid manipulation Gottesman (1997) under the action of the Clifford group. Despite their classical tractability, they can have high degrees of quantum entanglement Fattal et al. (2004).
@@
We start in Section II with a brief summary of the stochastic series expansion and introduce the idea of stabilizer states and their compact description. In Section III we demonstrate the use of stabilizers states in Quantum Monte Carlo simulations by picking a toy Hamiltonian and illustrating the algorithm in detail. Finally, Section IV presents a simulation of the transverse-field Ising model using this method and sketches how it may be extended to certain $\mathbb{Z}_{2}$ gauge theories. The GitHub repository Jeevanesan contains the C++ implementations of the stabilizer Monte Carlo algorithm and exact diagonalization programs used in this paper.
@@ II A brief review of SSE Quantum Monte Carlo and the Stabilizer Formalism
Consider a Hamiltonian $H$ on $N$ qubits. In a first step, the partition function is expanded in powers of $\beta$
$Z(\beta)$ $=$ $\text{Tr}[e^{-\beta H}]=\sum_{s}\langle s|e^{-\beta H}|s\rangle$
$=$ $\sum_{s}\sum_{n=0}^{\infty}\frac{(-\beta)^{n}}{n!}\langle s|H^{n}|s\rangle$
@@
$Z=\sum_{s}\sum_{i_{1},\dots,i_{n}}\sum_{n=0}^{\infty}\frac{(-\beta)^{n}}{n!}\langle s|H_{i_{1}}H_{i_{2}}\dots H_{i_{n}}|s\rangle$ (2)
@@
where the factor (${{L}\choose{L-n}}^{-1}$ compensates for overcounting identical terms. The sum is extended over all combinations of length-$L$ operator strings and $n$ refers to the number of non-identity operators in it. In other words, $L-n$ of the $H_{i}$ are now identity operators. In the second line, we introduced the notation $W$ for the weight of each configuration.
@@
To summarize, the partition function (II) has been transformed into a sum of matrix elements involving $L$ operators. Each term can be interpreted as the weight of a physical process where the system begins and ends in state $|s\rangle$ after the action of $L$ operators. Thus, if the Hamiltonian $H$ describes a lattice system of dimension $d$ then the representation (4) describes a system of dimension $d+1$ where the additional dimension is periodic with $L$ lattice sites Suzuki (1976).
@@
The crucial step in SSE Monte Carlo is that the decomposition of $H$ into terms $\sum_{i}H_{i}$ is done in such a way that the action of any $H_{i}$ on a computational basis state $|s\rangle$ yields another computational basis state $|s^{\prime}\rangle$. This can usually be accomplished by breaking $H$ into sufficiently small pieces and is called the no-branching provision. It guarantees that no superpositions over basis states are formed during the simulation process and thereby ensures easy bookkeeping. General superpositions of basis states are practically impossible to keep track of on a classical computer, except when system sizes are small.
@@
The sampling of the partition function now proceeds by setting up a Markov chain Metropolis et al. (1953); Diaconis (2009) with each element of the chain being a state $|s\rangle$ together with a string of operators. The weight of the Markov chain element is given by the corresponding term in eq. (3). An efficient simulation is only possible if the weights are all non-negative. Occasionally, one can manipulate the Hamiltonian terms $H_{i}$ by the addition of constants and render all weights non-negative. If this is not possible, one faces the Monte Carlo sign-problem. It can be dealt with, in a brute force manner, by absorbing the sign of the weight into the observable. But generally, this leads to long convergence times when calculating observable averages, and the variances diverge exponentially with system-size.
@@
A Markov chain update occurs by proposing a change of the state $|s\rangle$ or the insertion/deletion of an operator in the string. The proposal is accepted or rejected to satisfy detailed balance. The latter involves the ratios of the weights, an example appears in Section III.2. Note that if only one operator is inserted or removed, most factors in the weights are unchanged and will therefore cancel out. This leads to simple expressions for the update rules. However, since the application of the $H_{i_{k}}$ in eq. (3) always yields basis states, randomly inserting or rejecting an operator will often result in a state with zero weight. Instead, the insertion and removal of operators have to be carefully crafted. Furthermore, local updates of the operator string correspond to local changes of the periodic word line. Thus the system cannot explore all winding number sectors. For certain Hamiltonians, smart algorithms have been discovered that insert/delete a large number of operators simultaneously in order to achieve fast changes of the sampled configurations Syljuåsen and Sandvik (2002); Sandvik (2003, 1999).
@@
In this paper, we ask whether the no-branching condition is essential or if there is an efficient way to track the branching of the states. To this end, we return to the expression in eq. (3) and consider an interpretation of the sequence of operators as a quantum circuit. This is possible if each $H_{i}$ is either proportional to a unitary or to a projection operator, with the latter describing the outcomes of partial projective measurements. This observation by itself would not be particularly useful since an arbitrary quantum circuit still takes an exponential time to simulate classically. It has been proposed to evaluate the matrix element in (3) by means of quantum computers, resulting in a quantum SSE algorithm Tan et al. (2022); see Jiang et al. (2024) for a helpful review.
However, in the absence of large-scale quantum computers, progress can still be made through classical computation if the unitary operator is a Clifford circuit. In this case, efficient classical simulation is possible, as formalized by the Gottesman-Knill theorem Gottesman (1998a).
@@
Thus, acting with $U$ on $|\psi\rangle$ is equivalent to conjugating all the generators by $U$.
For general $U$, the resulting generators will lie outside the Pauli group and their description will be just as cumbersome as describing $|\psi\rangle$ in the computational basis. However, if $U$ maps all Pauli group elements to (possibly different) Pauli group elements under conjugation, then an update according to (9) is straightforward. Since conjugation by $U$ followed by conjugation by $V$ is the same as conjugation by $VU$, these elements form a group, the Clifford group. It was shown in Gottesman (1998b) that the Clifford group is finitely generated by the Hadamard-, phase- and controlled-not-gates. This set of gates is, of course, not universal. In fact, circuits made up of only these elements are simulable with classical algorithms in polynomial time. Adding $T$-unitaries to this gate set would make the circuit universal and generally intractable with classical simulations.
@@ III A Controlled-Not Toy Model
The two sums do not commute with each other. The CX term is a controlled-not operation with qubit $i$ being the control and qubit $i+1$ the target. In the Z-basis, this operator flips qubit $i+1$, conditioned on the state of qubit $i$ being $|1\rangle$. The $H^{(2)}$ term represents an external-field acting on each qubit individually. We work throughout this paper with periodic boundary conditions such that site $N+1$ is identified with site $1$. We also measure energy in units of $J$, i.e. we set $J=1$. Conditional interactions, similar to the CX operator, also appear in constrained statistical mechanics systems and lead to interesting dynamical effects. A recent example is the PXP model that describes Rydberg blockade physics Lesanovsky and Katsura (2012): an atom can only be excited if the neighboring atoms are in the ground state. The system shows interesting many-body dynamics Bernien et al. (2017) and quantum scarring Serbyn et al. (2021) that can be traced back to the constrained dynamics.
The Hamiltonian in eq. (10) is physically sensible because the unitary CX is also hermitian. Thus, in terms of Pauli operators, the Hamiltonian eq. (10) reads
@@
This simple example already illustrates how the consecutive application of the sequence of operators in eq. (15) results in superpositions of computational basis states with non-zero entanglement. The fact that the Bell-state appears in eq. (17) also demonstrates that in general these superpositions cannot be rewritten as product states through a local basis change. Nevertheless, the stabilizer formalism allows us to keep track of such states.
@@ III.1 Calculation of $\mathcal{M}$
$G_{m}$$\equiv$ $\gamma_{m}X_{1}^{P_{m1}}Z_{1}^{Q_{m1}}\otimes\dots\otimes X_{N}^{P_{mN}}Z_{N}^{Q_{mN}}$ (19)
$=$ $\gamma_{m}\prod_{n=1}^{N}X_{n}^{P_{mn}}Z_{n}^{Q_{mn}}$
@@
Thus we let $|\psi\rangle$ and $|\psi^{\prime}\rangle$ be states normalized to $1$ and let $F$ and $F^{\prime}$ keep track of the actual normalization. This step is necessary because the stabilizers do not determine the norm of a state.
After projector $\Pi_{i}$ acts on state $|\psi\rangle$, the new stabilizers can be worked out as follows. Since the generators are products of Pauli operators, the $X_{i}$ operator will either commute or anti-commute with them. Retain all the generators that commute with $X_{i}$. If all generators commute with $X_{i}$, we are done and set $F^{\prime}=F$.
@@
$\gamma^{\prime}_{m_{a}}$ $=$ $(-1)^{\rho}\gamma_{m_{1}}\gamma_{m_{a}}$ (31)
@@
The sign $(-1)^{\rho}$ stems from moving the $X$ operators in $G_{m_{a}}$ past the $Z$ operators in $G_{m_{1}}$. All other components of $\gamma$, $P$ and $Q$ are left unmodified.
@@
We first observe that $\langle s|s^{\prime}\rangle$ is non-negative. The reason for this is that both $\Pi_{i}$ and $[\text{CX}]_{i,i+1}$ are non-negative matrices in the $Z$ basis, i.e. all the entries are non-negative. The product of non-negative matrices is non-negative in that same basis, thus $\langle s|s^{\prime}\rangle$ is non-negative. This is necessary below in order to avoid the sign-problem.